### v2

Abstract
We consider a sender–receiver game in which the receiver’s action is binary and the sender’s preferences are state-independent. The state is multidimensional. The receiver can select one dimension of the state to check (i.e., observe) before choosing his action. We identify a class of influential equilibria in which the sender’s message reveals which components of the state are highest, and the receiver selects one of these components to check. The sender can benefit from communication if and only if she prefers one of these equilibria to the no-communication outcome. Similar equilibria exist when the receiver can check multiple dimensions.
Keywords: cheap talk, partial verification
JEL Classification Codes: D82, D83
1 Introduction
Less informed agents often turn to biased experts for guidance. Since Crawford and Sobel (1982), a large theoretical literature has demonstrated how cheap talk can influence a receiver’s action, even when there is a conflict of interest between the sender and the receiver. Chakraborty and Harbaugh (2010) and Lipnowski and Ravid (2020) show that cheap talk can be influential even when the sender has state-independent preferences. However, there remains an important case in which cheap talk cannot be influential: if the receiver’s action is binary and the sender strictly prefers the same action in every state. In this case, the receiver must choose the same (mixed) action after every equilibrium message. Otherwise, the sender would deviate by always sending the message that induces her preferred action with the highest probability.
There are many applications in which the sender prefers the same action in every state. For example, a salesperson wants a shopper to buy her product, no matter the product’s true quality. A job candidate wants to be hired, regardless of her skills. A politician wants a voter’s support, regardless of whether her platform would benefit the voter. Communication in these settings is commonly observed, but it cannot be explained by the standard theory of cheap talk communication.
In this paper, we show that costless communication can be influential in such settings if the receiver can gather additional information about the state after he receives the sender’s message. For instance, the shopper can inspect some attributes of the product; the employer can test the candidate on certain tasks; and the voter can research some aspects of the politician’s platform. Even though the receiver understands that the sender’s message is intended to maximize the probability of the sender’s preferred action, the sender’s message can influence which information the receiver acquires. The salesperson can guide the shopper to inspect the best attributes of the product; the job candidate can steer the recruiter to test her strongest skills; and the politician can encourage the voter to research the most compelling parts of her platform.
We consider a sender–receiver game, phrased in terms of our leading application of a seller and a buyer. The state (of some product) is a vector of $N$ binary attributes. The state is drawn from a common prior that is symmetric across the attributes. The receiver (he) chooses whether to buy the product (at a fixed, exogenous price). The sender (she) always strictly prefers the receiver to buy. The timing is as follows. The sender observes the state and then sends a message to the receiver. After seeing the message, the receiver chooses which attribute to “check.” The receiver observes whether the checked attribute is good or bad, and then he decides whether to buy the product. We interpret the restriction on verification as a time or cognitive constraint, as in Glazer and Rubinstein (2004). For instance, a shopper may not be able to learn about every technical specification of a smartphone.
We construct a family of influential equilibria, parameterized by $k=1,\ldots,N-1$. In a top-$k$ equilibrium, the sender’s message reveals which $k$ attributes are best, with ties broken by uniform randomization. Crucially, the message does not reveal the quality of these $k$ attributes or the ordering among these $k$ attributes. The receiver randomly selects one of these $k$ attributes to check. Then he buys if and only if that attribute is good. In these equilibria, the sender’s message induces the receiver to check an above-average attribute. In this way, communication can strictly benefit the sender by shaping the information that the receiver acquires.
Theorem 1 characterizes whether each top-$k$ equilibrium exists. This family of equilibria has an intuitive structure. As $k$ increases, the checked attribute becomes less upwardly biased. Therefore, the probability of purchase decreases, but the equilibrium can be sustained at higher prices.
The top equilibria are extremal within the set of all equilibria, in a sense that we make precise using Lemma 1. Therefore, by analyzing the top equilibria, we can draw conclusions about all equilibria. Theorem 2 shows that the sender can strictly benefit from communication if and only if the sender strictly prefers one of the top equilibria to the no-communication outcome. The key condition is that the top-$(N-1)$ equilibrium exists. Theorem 3 shows that the sender-preferred top-$k$ equilibrium is Pareto efficient within the set of all equilibria. Moreover, for each $k$, there is some price at which the top-$k$ equilibrium is sender-optimal among all equilibria.
Our main model assumes that the price is exogenous. Next, we allow the sender to choose the price (before observing the state). In this case, Theorem 4 shows that the sender-optimal outcome is either one of the top equilibria or else the no-communication outcome.
Finally, we consider an extension in which the receiver can check $n$ of the $N$ attributes, where $1<n<N$. We construct a natural analogue of each top-$k$ equilibrium. The sender’s message reveals which $k$ attributes are best, and the receiver checks attributes in that set. In Theorem 5, we characterize the interval of prices at which the top-$k$ equilibrium exists. As $k$ increases, this interval shifts upward.
The rest of the paper is organized as follows. After discussing related literature, we present the model in Section 2. In Section 3, we characterize whether the top equilibria exist. In Section 4, we analyze the optimality of the top equilibria, and we characterize whether the sender benefits from communication. In Section 5, we consider the extension in which the receiver can check multiple attributes. Section 6 is the conclusion. Proofs are in Appendix A.
Related literature
Since Crawford and Sobel (1982), the cheap talk literature has demonstrated the possibility of influential communication between players with partially aligned preferences. If the state and action spaces are multidimensional, and the players have state-dependent preferences, then the sender can credibly reveal the component of the state along a direction of agreement (Battaglini, 2002; Ambrus and Takahashi, 2008) or the ranking of the components of the state (Chakraborty and Harbaugh, 2007). More recent work considers the extreme case in which the sender’s preferences are state-independent. In Chakraborty and Harbaugh (2010), the receiver matches his action with his updated expectation of the state. If the sender’s utility is strictly quasiconvex in the receiver’s action, then the sender can strictly benefit by making comparative statements about different components of the state. In a more general model, Lipnowski and Ravid (2020) build upon Aumann and Hart (2003) to show that the sender’s maximal equilibrium payoff is the quasiconcave envelope of her value function in belief space.
The verification structure in our model follows Glazer and Rubinstein (2004). In that paper, the receiver can check one component of the multidimensional state before choosing a binary action. [^1] Glazer and Rubinstein (2006) study a complementary problem where the sender chooses which hard evidence to present to the receiver. In a more general evidence setting, Hart et al. (2017) establish an equivalence between equilibria (without commitment) and receiver-optimal mechanisms (with commitment). The sender has a strict, state-independent preference for one action over the other. Crucially, the receiver has commitment power. Their paper solves for the receiver-optimal mechanism. Then they show that the receiver-optimal mechanism can be supported as an equilibrium, without receiver commitment. By contrast, we study the sender’s gains from communication, without receiver commitment. Carroll and Egorov (2019) extend the setting of Glazer and Rubinstein (2004) by allowing the sender to have a general utility function over the receiver’s induced beliefs. They give a necessary and sufficient condition on the utility function for the receiver to perfectly learn the state. Their condition is not satisfied in our setting.
In many other models of verification, the receiver observes an exogenous signal about the state. [^2] An exception is Hancart (2024). The receiver chooses from an arbitrary set of Blackwell experiments about the state. That paper characterizes whether it is optimal for the receiver to choose a Blackwell-dominated experiment. Kattwinkel (2019) and Silva (2024) give conditions under which the receiver, with commitment power, can benefit from communication, i.e., screening the sender. Weksler and Zik (2025) give conditions under which communication can influence the receiver’s action, when the sender has private information about the informativeness of the receiver’s signal. In these papers, the sender’s message influences the sensitivity of the receiver’s action to his signal realization. In our model, by contrast, the sender’s message influences which signal the receiver observes. Finally, communication equilibria have been analyzed in models where the receiver can (a) detect lies with an exogenous probability (Balbuzanov, 2019; Dziuda and Salas, 2018), (b) choose to detect lies at a cost (Sadakane and Tam, 2023), or (c) choose to learn the state at a cost (Bijkerk et al., 2018; Venkatesh, 2024).
In our model, the receiver can verify the state, but the sender’s message is costless and unrestricted. An alternative modeling approach is taken in disclosure games (Milgrom, 1981; Grossman, 1981). In that model, the sender can “prove” to the receiver any true statement about the state. Thus, messages have intrinsic meaning, and the true state determines which messages are feasible. Subsequent work imposes structure on the class of true statements that can be proven. In Fishman and Hagerty (1990), the sender can disclose one dimension of a multidimensional state. In Dziuda (2011), the sender can present verifiable “arguments,” and the number of arguments is unknown to the receiver.
2 Model
There are two players: a sender (she) and a receiver (he). The sender privately observes the state $\theta\in\Theta=\{0,1\}^{N}$. Assume $N\geq 2$. The state is drawn from a full-support prior $\pi\in\Delta(\Theta)$ that is exchangeable (i.e., invariant to permutations of the $N$ dimensions). The receiver chooses a binary action $a\in\{0,1\}$. In our leading interpretation, the sender is the seller of a product and the receiver is the buyer. The action $a=1$ denotes buying. Each component $\theta_{i}$ represents a binary product attribute. [^3] Alternatively, the components can be interpreted as the outcomes of binary product tests (pass or fail), which are independently and identically distributed conditional on the (unobserved) product quality. The utilities of the sender and receiver are given by
$u_{S}(a,\theta)=ap,\qquad u_{R}(a,\theta)=a(v(\theta)-p),$
where $p>0$ is a fixed, exogenous price and $v(\theta)$ is the receiver’s consumption utility from the product in state $\theta$. We assume that the utility function $v\colon\Theta\to\mathbf{R}_{+}$ is strictly increasing and symmetric across the components. That is, $v$ can be expressed as a strictly increasing function of the number of good attributes, which we denote by $\lvert\theta\rvert\coloneqq\theta_{1}+\cdots+\theta_{N}$. The sender’s utility is equal to revenue, so the sender always wants the receiver to buy.
The receiver has a limited capacity to learn about the product. Following Glazer and Rubinstein (2004), we assume that the receiver can check exactly one component of the state $\theta$ [^4] In Section 5, we consider the case in which the receiver can check multiple components. Checking component $i$ perfectly reveals $\theta_{i}$. The timing is as follows. The sender observes the state and sends a costless message to the receiver. The receiver sees the message and then chooses which of the $N$ dimensions to check. After observing whether the checked component is good (11 11) or bad (00 0), the receiver chooses whether to buy the product.
Next, we define strategies. Let $M$ be a (sufficiently rich) message space. [^5] If $M$ is infinite, then we assume that $M$ is endowed with a $\sigma$-algebra and that all maps are measurable with respect to this $\sigma$-algebra. A messaging strategy for the sender is a function $m\colon\Theta\to\Delta(M)$, which assigns to each state a distribution over messages in $M$. Let $[N]=\{1,\ldots,N\}$. A strategy for the receiver is a pair $(c,b)$ consisting of (i) a checking strategy $c\colon M\to\Delta([N])$, which assigns to each message a probability distribution over attributes, and (ii) a buying strategy
$b\colon M\times[N]\times\{0,1\}\to[0,1],$
which specifies the probability that the receiver buys ($a=1$) as a function of the message, the attribute that is checked, and the value of that checked attribute. The solution concept is Nash equilibrium, which we simply call equilibrium. As in other cheap talk games, any Nash equilibrium outcome can be supported as a perfect Bayesian equilibrium (or even a sequential equilibrium).
3 Constructing the top equilibria
In this section, we construct a family of influential equilibria, dubbed the top equilibria. We characterize whether these equilibria exist.
3.1 Benchmark: No verification
As a benchmark, suppose that the receiver cannot check any component of the state, as in a standard cheap talk game. Since the sender always wants the receiver to buy, cheap talk cannot influence the receiver’s action. Indeed, the sender’s incentive constraints imply that all equilibrium messages must induce the same buying probability. If $p<\operatorname{\mathbf{E}}[v(\theta)]$, then in every equilibrium, the receiver buys the product. If $p>\operatorname{\mathbf{E}}[v(\theta)]$, then in every equilibrium, the receiver does not buy the product. If $p=\operatorname{\mathbf{E}}[v(\theta)]$, then the receiver is indifferent between buying and not buying. For each $q\in[0,1]$, there is an equilibrium in which the receiver buys with probability $q$.
3.2 Top equilibria
We now return to the main model with verification. We first consider what happens if the sender does not communicate with the receiver. Since the prior is exchangeable, the receiver is indifferent between checking each of the $N$ attributes. Whichever attribute he checks, he observes whether that attribute is bad or good, so his updated expectation about his consumption utility is either
$\underaccent{\bar}{\nu}\coloneqq\operatorname{\mathbf{E}}[v(\theta)\mid\theta_{1}=0]\quad\text{or}\quad\bar{\nu}\coloneqq\operatorname{\mathbf{E}}[v(\theta)\mid\theta_{1}=1].$
Clearly, $\underaccent{\bar}{\nu}<\bar{\nu}$. If $\underaccent{\bar}{\nu}<p<\bar{\nu}$, then the receiver buys the product if and only if the checked attribute is good. Of course, this no-communication outcome can be supported as a babbling equilibrium.
We now construct a family of influential equilibria, indexed by $k=1,\ldots,N-1$. In the top-$k$ equilibrium, the sender’s message reveals which $k$ attributes are best, with ties broken uniformly. The receiver randomly selects one of these $k$ “recommended” attributes to check. Then the receiver buys if and only if the checked attribute is good. For example, suppose that $N=3$. Consider the top-22 22 equilibrium. If $\theta=(1,0,0)$, then with probability $1/2$ the sender’s message indicates that “attributes 1 and 2 are the two best attributes,” and with probability $1/2$ the sender’s message indicates that “attributes 1 and 3 are the two best attributes.” To be sure, messages have no intrinsic meaning; their interpretation is determined endogenously by the sender’s messaging strategy.
Formally, for $k=1,\ldots,N-1$, the top-$k$ strategy profile $(m_{k};c_{k},b_{k})$ is defined as follows. Let $\mathcal{P}_{k}$ denote the collection of all $k$-element subsets of $[N]$. For each $A\in\mathcal{P}_{k}$, let $\lvert\theta_{A}\rvert=\textstyle\sum_{i\in A}\theta_{i}$. That is, $\lvert\theta_{A}\rvert$ counts the number of attributes in $A$ that are good. Set $M=\mathcal{P}_{k}$. [^6] Technically, the set $M$ is fixed. We identify $\mathcal{P}_{k}$ with some subset of $M$. We assume that the receiver treats every message $m\in M\setminus\mathcal{P}_{k}$ as if he had received the same fixed message $m_{0}\in\mathcal{P}_{k}$. For each state $\theta$, let $m_{k}(\theta)$ be the uniform distribution over $\operatorname*{argmax}\nolimits_{A\in\mathcal{P}_{k}}\,\lvert\theta_{A}\rvert$. For each $A\in\mathcal{P}_{k}$, let $c_{k}(A)$ be the uniform distribution over $A$. Finally, let $b_{k}(A,i,z)=z$ for each $(A,i,z)\in\mathcal{P}_{k}\times[N]\times\{0,1\}$. That is, the receiver buys if and only if the attribute that he checks is good. If the top-$k$ strategy profile is an equilibrium, then we call it the top-$k$ equilibrium, and we say that the top-$k$ equilibrium exists.
Our first result gives a necessary and sufficient condition for the existence of the top-$k$ equilibrium. To state this condition, we introduce notation for the signal that the receiver obtains under the top-$k$ strategy profile. For each $k\in[N]$, let $T^{k}$ ($B^{k}$) denote a uniform draw from the top (bottom) $k$ attributes, where these top (bottom) attributes are determined by uniform tie-breaking. In state $\theta$, we know that among the top $k$ attributes, $\lvert\theta\rvert\wedge k$ are good; among the bottom $k$ attributes, $(\lvert\theta\rvert-(N-k))_{+}$ are good. Therefore, given the state $\theta$, the random variables $T^{k}$ and $B^{k}$ are conditionally independent Bernoulli random variables, with
$\mathbf{P}(T^{k}=1\mid\theta)=\frac{\lvert\theta\rvert\wedge k}{k},\qquad\mathbf{P}(B^{k}=1\mid\theta)=\frac{(\lvert\theta\rvert-(N-k))_{+}}{k}.$
For $k=1,\ldots,N-1$, define the thresholds
$\underaccent{\bar}{p}_{k}=\operatorname{\mathbf{E}}[v(\theta)\mid T^{k}=0],\qquad\bar{p}_{k}=\operatorname{\mathbf{E}}[v(\theta)\mid(T^{k},B^{N-k})=(1,0)].$
We will interpret these thresholds after stating the equilibrium condition.
Theorem 1 (Top-$k$ equilibrium condition)
For $k=1,\ldots,N-1$, the top-$k$ equilibrium exists if and only if
$\underaccent{\bar}{p}_{k}\leq p\leq\bar{p}_{k}.$
Moreover, these price thresholds are ordered as follows:
$v(\bm{0})=\underaccent{\bar}{p}_{1}<\cdots<\underaccent{\bar}{p}_{N-1}<\underaccent{\bar}{\nu}<\bar{p}_{1}<\cdots<\bar{p}_{N-1}<\bar{\nu}.$
For each $k$, the top-$k$ equilibrium exists if and only if the price $p$ lies in the nondegenerate interval $[\underaccent{\bar}{p}_{k},\bar{p}_{k}]$. The lower bound $p\geq\underaccent{\bar}{p}_{k}$ ensures that if the receiver checks one of the $k$ recommended attributes and sees that it is bad, then he finds it optimal to not buy. The upper bound $p\leq\bar{p}_{k}$ ensures that the receiver finds it optimal to check one of the recommended attributes, rather than one of the unrecommended attributes. Suppose that the receiver deviates by checking one of the unrecommended attributes and then buying if and only if that attribute is good. Under this deviation, the receiver’s buying decision differs from the equilibrium in the event that the checked unrecommended attribute is bad, but the recommended attribute that the receiver would have checked is good. The inequality $p\leq\bar{p}_{k}$ ensures that the receiver weakly prefers buying to not buying, conditional on this event, which corresponds to $(B^{N-k},T^{k})=(0,1)$.
Figure 1 illustrates these price thresholds and the associated buying probabilities in a numerical example. In this example, $N=4$ and $v(\theta)=\lvert\theta\rvert$. The state distribution is such that $\lvert\theta\rvert$ is equally likely to take the values $0,\ldots,4$. For each $k=1,2,3$, we plot the buying probability $\mathbf{P}(T^{k}=1)$ over the range of prices at which the top-$k$ equilibrium exists. The price range $[\underaccent{\bar}{p}_{k},\bar{p}_{k}]$ shifts up as $k$ increases. Intuitively, as $k$ increases, the recommended attributes become less upwardly biased relative to a randomly chosen attribute. The dashed line indicates the buying probability without communication: for $p<\underaccent{\bar}{\nu}$, the receiver always buys; for $\underaccent{\bar}{\nu}<p<\bar{\nu}$, the receiver buys if and only if the checked attribute is good. [^7] For $p>\bar{\nu}$, the receiver never buys. This part of the graph is omitted from the plot. If $\underaccent{\bar}{\nu}<p\leq\bar{p}_{3}$, then there exists a top equilibrium that the sender strictly prefers to the no-communication outcome.
[…]
Figure 1: Top equilibrium conditions and associated buying probabilities
4 Equilibrium welfare
In this section, we use the top equilibrium construction to characterize whether the sender can strictly benefit from communication. We also give conditions under which each top equilibrium is Pareto efficient or sender-optimal among all equilibria.
4.1 Equilibrium bounds
Under each strategy profile, the expected payoffs of the sender and the receiver are pinned down by the buying probabilities conditional on each realization of $\lvert\theta\rvert$. For each $j$, let $q_{j}$ denote the buying probability conditional on the event that $\lvert\theta\rvert=j$. We use the sender’s incentive constraints to derive bounds on these buying probabilities that are satisfied by every equilibrium.
Lemma 1 (Equilibrium buying probabilities)
In any equilibrium, the induced buying probabilities $q_{0},\ldots,q_{N}$ satisfy
$q_{j-1}\geq\frac{j-1}{j}q_{j},\qquad j=1,\ldots,N.$ (1)
Lemma 1 uses only the sender’s incentive constraints, so the bound in (1) is independent of the price $p$ and the receiver’s utility function $v$. The sender’s incentive constraints limit how steeply the buying probability can increase as a function of the number $\lvert\theta\rvert$ of good attributes. The precision of the verification technology determines the maximal steepness. Indeed, without verification, the buying probability must be flat. If the receiver could perfectly verify the entire state, then the buying probability could be arbitrarily steep.
To prove Lemma 1, we fix an arbitrary equilibrium and then analyze a special class of deviations. Consider some type $\theta$ with $\lvert\theta\rvert=j>1$. Choose a dimension $i$ such that $\theta_{i}=1$, and consider the state $\theta^{\prime}=\theta-e_{i}$, where $e_{i}$ is the $i$-th standard basis vector. Type $\theta^{\prime}$ can mimic type $\theta$ by sending some message that is sent by type $\theta$ in equilibrium. If the receiver checks any dimension other than dimension $i$, then he cannot distinguish type $\theta$ from type $\theta^{\prime}$, so the outcome must be the same as if type $\theta$ had sent that message. There are $j$ feasible choices for this dimension $i$. Averaging over these $j$ choices, we conclude that the average buying probability over those $j$ states is at least $(j-1)/j$ times the buying probability in state $\theta$. By further averaging over all states $\theta$ with $j$ good attributes, we obtain the inequality in (1).
For $k=1,\ldots,N-1$, let $q^{k}=(q_{0}^{k},\ldots,q_{N}^{k})$ denote the vector of buying probabilities induced by the top-$k$ strategy profile. We have
$q_{j}^{k}=\mathbf{P}(T^{k}=1\mid\lvert\theta\rvert=j)=\frac{j\wedge k}{k},\qquad j=0,\ldots,N.$ (2)
The vector $q^{k}$ satisfies (1) with equality for each $j$ such that $q_{j-1}^{k}<1$. Let $Q$ be the set of all vectors $q$ in $[0,1]^{N+1}$ that satisfy (1). The set $Q$ is a polytope. It is easily verified that each vector $q^{k}$ is an extreme point of $Q$. Let $Q_{0}$ be the set of all buying probability vectors that are induced by some equilibrium. The set $Q_{0}$, unlike $Q$, depends on the price $p$ and the utility function $v$. This set $Q_{0}$ is difficult to characterize, but we know from Lemma 1 that $Q_{0}$ is a subset of $Q$. It follows that the vector $q^{k}$ is an extreme point of $Q_{0}$ whenever the top-$k$ equilibrium exists.
4.2 Benefiting from communication
We say that the sender benefits from communication if there exists an equilibrium that the sender strictly prefers to the (sender-preferred) no-communication outcome.
Theorem 2 (Benefiting from communication)
The sender benefits from communication if and only if
$\underaccent{\bar}{\nu}<p\leq\bar{p}_{N-1}.$
Thus, the sender benefits from communication if and only if there exists a top equilibrium that the sender strictly prefers to the (sender-preferred) no-communication outcome. The price range in Theorem 2 can be visualized in Figure 1, where $N-1=3$.
We argue by cases. If $p\leq\underaccent{\bar}{\nu}$, then without communication, the receiver is willing to always buy the good. This is the sender’s first-best outcome, so the sender cannot benefit from communication.
Next, suppose $\underaccent{\bar}{\nu}<p\leq\bar{\nu}$. Over this range, without communication, the receiver buys if and only if the checked attribute is good. Recall from Theorem 1 that $\underaccent{\bar}{\nu}<\bar{p}_{N-1}<\bar{\nu}$. If $p\leq\bar{p}_{N-1}$, then the top-$(N-1)$ equilibrium exists. Thus, the sender strictly benefits from communication by guiding the receiver to check attributes that are better than average. If $p>\bar{p}_{N-1}$, then no top equilibrium exists. In principle, there could be other equilibria that the sender prefers to the no-communication outcome. Our proof rules this out, using the bounds in Lemma 1.
Finally, if $p>\bar{\nu}$, then without communication, the receiver never buys. In this case, we use Lemma 1 to show that there is no equilibrium in which the receiver buys with positive probability. The sender’s incentive constraints are critical to this conclusion. For prices $p$ slightly above $\bar{\nu}$, the receiver would be willing to buy upon seeing that one of the worst attributes was good. But in equilibrium, the sender cannot induce the receiver to check a below-average attribute. The sender would strictly prefer the receiver, while thinking he was checking one of the worst attributes, to actually check one of the best.
4.3 Optimality of top equilibria
We say that an equilibrium is Pareto efficient if it is not Pareto dominated by any other equilibrium. An equilibrium is sender-optimal if no other equilibrium gives the sender a strictly higher payoff.
Theorem 3 (Optimality of top equilibria)
For $k=1,\ldots,N-1$, the top-$k$ equilibrium is Pareto efficient if $\bar{p}_{k-1}\leq p\leq\bar{p}_{k}$ and sender-optimal if $p=\bar{p}_{k}$.
By Theorem 1, the top-$k$ equilibrium is the sender-preferred top equilibrium if and only if $\bar{p}_{k-1}\leq p\leq\bar{p}_{k}$. Therefore, Theorem 3 implies that the sender-preferred top equilibrium is Pareto efficient. To prove Theorem 3, we show that if $\bar{p}_{k-1}\leq p\leq\bar{p}_{k}$, then no vector of buying probabilities that satisfies the bounds in Lemma 1 can Pareto dominate the top-$k$ equilibrium. If $p=\bar{p}_{k}$, then sender optimality follows from Pareto efficiency because the top-$k$ equilibrium gives the receiver the lowest possible payoff among all equilibria—the receiver’s payoff from the top-$k$ equilibrium is the same as his payoff from checking a random attribute and buying if and only if it is good. The prices $\bar{p}_{1},\ldots,\bar{p}_{N-1}$ can be visualized in Figure 1 as the right endpoints of the equilibrium price intervals.
Up to this point, we have assumed that the price $p$ is exogenous. Next, we assume that the sender can choose the price (before observing the state). Thus, the sender can choose the price to maximize her revenue in the resulting equilibrium. We characterize which prices can be optimal. To state this result, it is convenient to extend our definition of the top-$k$ equilibrium to the edge cases $k=0$ and $k=N$. Under the top-00 0 strategy profile, the sender conveys no information; the receiver checks an arbitrary attribute and then always buys. Under the top-$N$ strategy profile, the sender conveys no information; the receiver checks an arbitrary attribute and buys if and only if the checked attribute is good. [^8] In each case, there are multiple strategy profiles with this property, but all such strategy profiles are payoff equivalent. Let $\bar{p}_{0}=\underaccent{\bar}{\nu}$ and $\bar{p}_{N}=\bar{\nu}$. For $k=0$ and $k=N$, it is clear that $\bar{p}_{k}$ is the highest price at which the top-$k$ strategy profile is an equilibrium.
Theorem 4 (Optimal price and equilibrium)
Suppose that the sender chooses the price (before observing the state). There exists $k^{\ast}\in\{0,\ldots,N\}$ such that the sender’s payoff is maximized by the price $p=\bar{p}_{k^{\ast}}$ and the top-$k^{\ast}$ equilibrium.
[…]
Figure 2: The set $\mathcal{X}$ and its upper envelope
We illustrate the intuition for Theorem 4 graphically. Recall that $Q$ denotes the set of buying probability vectors $q=(q_{0},\ldots,q_{N})\in[0,1]^{N+1}$ satisfying the bounds in Lemma 1. The sender’s and receiver’s payoffs from a vector $q\in Q$ are determined by two statistics: the ex-ante buying probability $x(q)$ and the expected consumption utility $y(q)$. In particular, the sender’s payoff is $px(q)$ and the receiver’s payoff is $y(q)-px(q)$. To compute these statistics, we introduce additional notation. For $j=0,\ldots,N$, let $\pi_{j}$ denote the probability $\mathbf{P}(\lvert\theta\rvert=j)$, and let $v_{j}$ denote the consumption utility $v(\theta)$ in any state $\theta$ with $\lvert\theta\rvert=j$ [^9] This is well-defined because $v$ is symmetric across the components. We have
$x(q)=\sum_{j=0}^{N}\pi_{j}q_{j}\quad\text{and}\quad y(q)=\sum_{j=0}^{N}\pi_{j}q_{j}v_{j}.$
Let $\mathcal{X}=\{(x(q),y(q)):q\in Q\}$. The set $\mathcal{X}$ is a polytope since it is the image of the polytope $Q$ under the linear transformation $q\mapsto(x(q),y(q))$. Note that $\mathcal{X}$ depends on the state distribution and the utility function $v$, but not on the price $p$.
The set $\mathcal{X}$ is more convenient to analyze than the set $Q$ because it lies in $\mathbf{R}^{2}$ rather than $\mathbf{R}^{N+1}$. Figure 2 plots the set $\mathcal{X}$ in the numerical example with $N=4$ used in Figure 1. The upper envelope of $\mathcal{X}$ is highlighted. The extreme points of this upper envelope correspond to the top-$k$ equilibria for $k=0,\ldots,N$ and also the origin, as we show in Lemma 3 (Section A.4). At price $p$, the receiver’s indifference curves in $\mathcal{X}$ are parallel lines with slope $p$. For $k=0,\ldots,N-1$, at price $\bar{p}_{k}$ the receiver is indifferent between the top-$k$ equilibrium and deviating by checking a random attribute and buying if and only if that checked attribute is good. Therefore, $\bar{p}_{k}$ is the slope of the line segment between top-$k$ equilibrium vector and the top-$N$ equilibrium vector. Using the characterization of the upper envelope of $\mathcal{X}$, we compute for each price $p$ an upper bound $V(p)$ on the sender’s revenue from all equilibria at price $p$. For $k=0,\ldots,N$, the bound $V(\bar{p}_{k})$ equals the sender’s revenue from the top-$k$ equilibrium at price $\bar{p}_{k}$. To prove Theorem 4, it suffices to show that the function $V$ is globally maximized at some price in $\{\bar{p}_{0},\ldots,\bar{p}_{N}\}$. The function $V$ has kinks at these prices. We show that $V$ is convex between consecutive kinks, and hence the maximum must be achieved at one of the kinks.
The optimal price depends on the state distribution and the utility function $v$. Consider the simple example with $N=2$ and $v(\theta)=\lvert\theta\rvert$. It can be verified that the top-11 11 equilibrium is optimal if $\pi_{1}\geq 2\pi_{2}$ and the top-22 22 equilibrium is optimal if $\pi_{1}\leq 2\pi_{2}$. For the calculations, see Section A.8. When $\lvert\theta\rvert=1$, the receiver buys with certainty in the top-11 11 equilibrium but only with probability $1/2$ in the top-22 22 equilibrium. Thus, as $\pi_{1}/\pi_{2}$ increases, the buying probability ratio between the top-11 11 equilibrium and the top-22 22 equilibrium increases. The price ratio $\bar{p}_{1}/\bar{p}_{2}$ also increases because the receiver, upon seeing that a randomly chosen attribute is 11 11, assigns lower probability to the event $\lvert\theta\rvert=2$.
5 Checking multiple attributes
In the baseline model, the receiver can check exactly one attribute. In this section, we assume that the receiver has the capacity to simultaneously check $n$ of the $N$ attributes, where $1<n<N$. We rule out the case of perfect verification $(n=N)$: if the receiver can learn the state perfectly, then there is no role for communication.
First, we define strategies in this extended model. As in the baseline model, a messaging strategy for the sender is a function $m\colon\Theta\to\Delta(M)$. Recall that $\mathcal{P}_{n}$ denotes the collection of all $n$-element subsets of $[N]$. A strategy for the receiver is a pair $(c,b)$, consisting of a checking strategy $c\colon M\to\Delta(\mathcal{P}_{n})$, which specifies which $n$ attributes are checked, and a buying strategy
$b\colon M\times\mathcal{P}_{n}\times\{0,1\}^{n}\to[0,1],$
which specifies the probability that the receiver buys ($a=1$) as a function of the message, the set of attributes that are checked, and the values of those attributes.
We now extend the notion of a top-$k$ equilibrium to this setting. The top-$k$ messaging strategy for the sender is defined as in the baseline model: the sender’s message indicates the top $k$ attributes, with ties broken uniformly. The receiver’s top-$k$ strategy is as follows. The receiver checks $n\wedge k$ of the recommended attributes. Then the receiver buys if and only if the number of good attributes among those checked is at least some threshold, which is between 11 11 and $n\wedge k$ [^10] Technically, we have assumed that the receiver must check $n$ attributes. Thus, the receiver can check $n-(n\wedge k)$ of the unrecommended attributes, but their values do not affect his buying probability. We say that the top-$k$ equilibrium exists if such a strategy profile is an equilibrium. [^11] In order to speak of the top-$k$ equilibrium, we assume that if there are multiple such thresholds that make the receiver’s strategy a best response, the smallest threshold is chosen.
As in the baseline model, we give a necessary and sufficient condition for the existence of the top-$k$ equilibrium. To state the condition, we generalize our notation. Let $T_{j}^{k}$ ($B_{j}^{k}$) denote the number of good attributes in a uniform sample of $j\wedge k$ of the top (bottom) $k$ attributes, where these top (bottom) attributes are determined by uniform tie-breaking. For $k=1,\ldots,N-1$, define the thresholds
$\underaccent{\bar}{p}_{k,n}=\operatorname{\mathbf{E}}[v(\theta)\mid T_{n}^{k}=0],\qquad\bar{p}_{k,n}=\operatorname{\mathbf{E}}[v(\theta)\mid(T_{n}^{k},B^{N-k}_{n})=(n\wedge k,0)].$
With $n=1$, these definitions coincide with the definitions from the main model: $\underaccent{\bar}{p}_{k,1}=\underaccent{\bar}{p}_{k}$ and $\bar{p}_{k,1}=\bar{p}_{k}$.
Theorem 5 (Top-$k$ equilibrium condition with multiple attributes checked)
Suppose $1<n<N$. For $k=1,\ldots,N-1$, the top-$k$ equilibrium exists if and only if
$\underaccent{\bar}{p}_{k,n}\leq p\leq\bar{p}_{k,n}.$
Moreover, $\underaccent{\bar}{p}_{k,n}<\bar{p}_{k,n}$ for each $k$, and the price thresholds are ordered as follows:
$\underaccent{\bar}{p}_{1,n}<\cdots<\underaccent{\bar}{p}_{N-1,n}\quad\text{and}\quad\bar{p}_{1,n}<\cdots<\bar{p}_{N-1,n}.$
For each fixed $n$ and $k$, the top-$k$ equilibrium exists if and only if the price $p$ lies in the nondegenerate interval $[\underaccent{\bar}{p}_{k,n},\bar{p}_{k,n}]$. This equilibrium price interval shifts up as $k$ increases, just as in the baseline model with $n=1$. Intuitively, as the sender recommends a greater share of the attributes, the sample of recommended attributes becomes less upwardly biased, so the equilibrium can be sustained at higher prices.
When the receiver can check multiple attributes, the set of deviations is much richer. To prove Theorem 5, the trick is to compute the receiver’s best response in a hypothetical specification in which the receiver checks $n\wedge k$ of the $k$ recommended attributes and $n\wedge(N-k)$ of the $N-k$ unrecommended attributes. For simplicity, suppose here that $n\leq k$ and $n\leq N-k$. The random vector $(T_{n}^{k},B_{n}^{N-k})$ can only take the values
$(0,0),(1,0),\ldots,(n,0),(n,1),\ldots,(n,n).$
The receiver’s best response is to buy if and only if the value of $(T_{n}^{k},B_{n}^{N-k})$ exceeds some threshold vector. In order to implement this best response, the receiver does not need to observe both $T_{n}^{k}$ and $B_{n}^{N-k}$. First, suppose that the threshold is $(t,0)$ for some $t\leq n$. Since $(T_{n}^{k},B_{n}^{N-k})\geq(t,0)$ if and only if $T_{n}^{k}\geq t$, the receiver can implement this best response by checking $n$ of the recommended attributes. Conversely, suppose that the threshold is $(n,b)$ for some $b>0$. Since $(T_{n}^{k},B_{n}^{N-k})\geq(n,b)$ if and only if $B_{n}^{N-k}\geq b$, the receiver can implement this best response by checking $n$ of the unrecommended attributes; the receiver cannot implement this best response by checking $n$ of the recommended attributes. The upper bound $p\leq\bar{p}_{k,n}$ ensures that the threshold is at most $(n,0)$, and hence it is optimal for the receiver to check the recommended attributes. The lower bound $p\geq\underaccent{\bar}{p}_{k,n}$ ensures that the threshold is strictly above $(0,0)$, so that after the receiver checks these recommended attributes, he does not always buy.
The price thresholds satisfy an intuitive order, as in the baseline model. The behavior of the buying probability is now more subtle. As $p$ and $k$ vary, so does the number of checked attributes that must be good in order for the receiver to buy.
Remark (Adaptive checking).
We have assumed that the receiver selects all $n$ attributes before seeing any of their values. If the receiver could instead check the attributes adaptively, based on the realizations of previously checked attributes, then our proof of Theorem 5 would still go through. Checking $n$ attributes adaptively is still less informative than checking $n\wedge k$ of the recommended attributes and $n\wedge(N-k)$ of the unrecommended attributes.
6 Conclusion
In this paper, we study how a sender with state-independent preferences can benefit from costless communication when the receiver can partially verify the state. We identify a natural class of equilibria, the top equilibria, in which the sender guides the receiver toward the best components of the state. Because these equilibria are extremal within the set of all equilibria, this construction allows us to draw conclusions about all equilibria. In particular, the sender can benefit from communication if and only if she prefers one of these equilibria to the no-communication outcome.
Following Glazer and Rubinstein (2004), we have assumed that checking attributes is costless up to some checking capacity, and prohibitively costly beyond that capacity. This captures a hard constraint on the receiver’s time or cognitive resources. Incorporating a more flexible model of information acquisition is an interesting direction for future work.
Appendix A Proofs
Throughout the proofs, we adopt the following notation. For each $j=0,\ldots N$, let $\pi_{j}=\mathbf{P}(\lvert\theta\rvert=j)$ and let $v_{j}$ denote the value of $v(\theta)$ when $\lvert\theta\rvert=j$.
A.1 Favorability
Given $n,k\in\{1,\ldots,N\}$, we consider the joint distribution of $(|\theta|,T_{n}^{k},B_{n}^{N-k})$. Note that the support of $(T_{n}^{k},B_{n}^{N-k})$ is
$S(n,k)\coloneqq\{(0,0),\ldots,(n\wedge k,0),(n\wedge k,1),\ldots,(n\wedge k,n\wedge(N-k))\}.$
This set is totally ordered by the product order.
We are interested in the distribution of $\lvert\theta\rvert$, conditional on events involving $(T_{n}^{k},B_{n}^{N-k})$. For any such event $E$, let
$\operatorname{supp}E=\{j\in\{0,\ldots,N\}:\mathbf{P}(E\mid\lvert\theta\rvert=j)>0\}.$
Note that this notion is different from the standard notion of the support of a random variable. Consider events $E$ and $E^{\prime}$ with positive probability. We say that $E$ is weakly more favorable than $E^{\prime}$ if the ratio
$\frac{\mathbf{P}(E\mid\lvert\theta\rvert=j)}{\mathbf{P}(E^{\prime}\mid\lvert\theta\rvert=j)}$ (3)
is weakly increasing in $j$ over $\operatorname{supp}E\cup\operatorname{supp}E^{\prime}$; we use the convention that $a/0=\infty$ for $a>0$. If, further, the ratio in (3) is not constant over $\operatorname{supp}E\cup\operatorname{supp}E^{\prime}$, then we say that $E$ is strictly more favorable than $E^{\prime}$. [^12] Our terminology is inspired by Milgrom (1981). Unlike in his definition, we do not assume that the events have full support. Our formal definition mirrors the definition of the likelihood ratio order in Shaked and Shanthikumar (2007, 1.C.1, p. 42). If $E$ is strictly more favorable than $E^{\prime}$, then $\operatorname{\mathbf{E}}[v(\theta)\mid E]>\operatorname{\mathbf{E}}[v(\theta)\mid E^{\prime}]$, by standard results about the likelihood ratio order.131313 If $E$ is weakly more favorable than $E^{\prime}$, then the distribution of $\lvert\theta\rvert$, conditional on $E$, likelihood ratio dominates the distribution of $\lvert\theta\rvert$, conditional on $E^{\prime}$, in the sense of Shaked and Shanthikumar (2007, 1.C.1, p. 42). Moreover, likelihood ratio dominance implies first-order stochastic dominance (Shaked and Shanthikumar, 2007, Theorem 1.C.1, p. 43).
A.2 Proof of Theorem 1
In Section A.9, we prove a more general version of Theorem 5 that also covers the case $n=1$. Moreover, we can apply Lemma 5 (Section A.9) with $k=N$ and $n=1$ to conclude that $\underaccent{\bar}{p}_{N-1,1}<\underaccent{\bar}{p}_{N,1}=\underaccent{\bar}{\nu}$ and $\bar{p}_{N-1,1}<\bar{p}_{N,1}=\bar{\nu}$. Thus, it remains to prove that $\underaccent{\bar}{\nu}<\bar{p}_{1}$. We prove that the event $(T^{1},B^{N-1})=(1,0)$ is strictly more favorable than the event $\theta_{1}=0$; see Section A.1 for the definition. The event $(T^{1},B^{N-1})=(1,0)$ has support $\{1,\ldots,N-1\}$. The event $\theta_{1}=0$ has support $\{0,\ldots,N-1\}$. We have
$\frac{\mathbf{P}\left((T^{1},B^{N-1})=(1,0)\mid\lvert\theta\rvert=0\right)}{\mathbf{P}(\theta_{1}=0\mid\lvert\theta\rvert=0)}=0,$
and for $j=1,\ldots,N-1$, we have
$\frac{\mathbf{P}\left((T^{1},B^{N-1})=(1,0)\mid\lvert\theta\rvert=j\right)}{\mathbf{P}(\theta_{1}=0\mid\lvert\theta\rvert=j)}=\frac{(N-j)/(N-1)}{(N-j)/N}=\frac{N}{N-1}.$
A.3 Proof of Lemma 1
Fix an equilibrium $(m;c,b)$. For each type $\theta$, choose a message $\bar{m}(\theta)$ that is sent by type $\theta$ in this equilibrium. [^14] Technically, if $M$ is infinite, choose a message $\bar{m}(\theta)$ in $M$ that maximizes the buying probability for type $\theta$, given the receiver’s strategy $(c,b)$. For each $i=1,\ldots,N$, let
$d_{i}^{0}(\theta)=c_{i}(\bar{m}(\theta))b(\bar{m}(\theta),i,0),\qquad d_{i}^{1}(\theta)=c_{i}(\bar{m}(\theta))b(\bar{m}(\theta),i,1).$
Therefore, in each state $\theta$, the equilibrium buying probability, $q(\theta)$, is given by
$q(\theta)=\sum_{i=1}^{N}\left[d_{i}^{0}(\theta)+\theta_{i}(d_{i}^{1}(\theta)-d_{i}^{0}(\theta))\right].$
For all $\theta,\theta^{\prime}\in\Theta$, type $\theta^{\prime}$ can mimic type $\theta$ by sending message $\bar{m}(\theta)$, so
$q(\theta^{\prime})\geq\sum_{i=1}^{N}\left[d_{i}^{0}(\theta)+\theta^{\prime}_{i}(d_{i}^{1}(\theta)-d_{i}^{0}(\theta))\right].$ (4)
Say that type $\theta^{\prime}$ directly precedes type $\theta$, denoted $\theta^{\prime}\prec\theta$, if $\theta=\theta^{\prime}+e_{i}$ for some index $i$, where $e_{i}$ is the $i$-th standard basis vector. Fix $j\in\{1,\ldots,N\}$. We may assume $j>1$ since the desired inequality (1) is trivial if $j=1$. For each fixed state $\theta$ with $\lvert\theta\rvert=j$, average (4) over the $j$ states $\theta^{\prime}$ that directly precede $\theta$. Since
$\frac{1}{j}\sum_{\theta^{\prime}:\theta^{\prime}\prec\theta}\theta_{i}^{\prime}=\frac{j-1}{j}\theta_{i},\qquad i=1,\ldots,N,$
we get
$\frac{1}{j}\sum_{\theta^{\prime}:\theta^{\prime}\prec\theta}q(\theta^{\prime})$ $\geq\sum_{i=1}^{N}\left[d_{i}^{0}(\theta)+\frac{j-1}{j}\theta_{i}(d_{i}^{1}(\theta)-d_{i}^{0}(\theta))\right]$ (5)
$\geq\frac{j-1}{j}\sum_{i=1}^{N}\left[d_{i}^{0}(\theta)+\theta_{i}(d_{i}^{1}(\theta)-d_{i}^{0}(\theta))\right]$
$=\frac{j-1}{j}q(\theta),$
where the second inequality holds because $d_{i}^{0}(\theta)\geq 0$ for each $i$. Now, average (5) over all $\tbinom{N}{j}$ states $\theta$ with $\lvert\theta\rvert=j$ to get
$q_{j-1}\geq\frac{j-1}{j}q_{j}.$
A.4 Preliminaries for Theorems 24
To prove the remaining theorems, we introduce additional notation. For $k=1,\ldots,N$, define $q^{k}=(q_{0}^{k},\ldots,q_{N}^{k})$ by
$q_{j}^{k}=\frac{j\wedge k}{k},\qquad j=0,\ldots,N.$
Thus, $q^{k}$ represents the buying probability vector induced by the top-$k$ equilibrium strategy. This extends the definition in (2) to the case $k=N$. To cover edge cases, we let $q^{0}=\bm{1}$ and $q^{N+1}=\bm{0}$. Observe that
$q^{N+1}\leq q^{N}\leq\cdots\leq q^{1}\leq q^{0}.$
For $k=1,\ldots,N$, define $\Delta^{k}=(\Delta_{0}^{k},\ldots,\Delta_{N}^{k})$ by
$\Delta_{j}^{k}=\frac{j}{k}[j\leq k],\qquad j=0,\ldots,N.$
Define $\Delta^{0}=(\Delta_{0}^{0},\ldots,\Delta_{N}^{0})$ by setting $\Delta_{0}^{0}=1$ and $\Delta_{j}^{0}=0$ for $j>0$.
Next, we define some special utility expressions. For $k=0,\ldots,N$, let
$u_{k}(p)=\sum_{j=0}^{N}\pi_{j}\Delta^{k}_{j}(v_{j}-p).$
For $k=0,\ldots,N-1$, we have $q^{k}-q^{k+1}=\Delta^{k}/(k+1)$, so
$\sum_{j=0}^{N}\pi_{j}(q_{j}^{k}-q_{j}^{k+1})(v_{j}-p)=\frac{u_{k}(p)}{k+1}.$ (6)
And for the case $k=N$, we have
$\sum_{j=0}^{N}\pi_{j}(q_{j}^{N}-q_{j}^{N+1})(v_{j}-p)=u_{N}(p).$ (7)
For the next result, recall that $\bar{p}_{0}=\underaccent{\bar}{\nu}$ and $\bar{p}_{N}=\bar{\nu}$.
Lemma 2 (Price thresholds)
The following hold.
  1. (i)
    For $k\in\{0,\ldots,N-1\}$, we have: $p\leq\bar{p}_{k}\iff\sum_{j=0}^{N}\pi_{j}(q^{k}_{j}-q^{N}_{j})(v_{j}-p)\geq 0$.
  2. (ii)
    For $k\in\{1,\ldots,N-2\}$, we have: $p\geq\bar{p}_{k}\implies u_{k}(p)<0$.
  3. (iii)
    For $k\in\{N-1,N\}$, we have: $p\leq\bar{p}_{k}\iff u_{k}(p)\geq 0$.
Recall that $\mathcal{X}=\{(x(q),y(q)):q\in Q\}$. For $x\in[0,1]$, let $y^{\ast}(x)=\max\{y:(x,y)\in\mathcal{X}\}$; this maximum exists because $\mathcal{X}$ is compact (and the projection of $\mathcal{X}$ onto the horizontal axis is $[0,1]$). The upper envelope of $\mathcal{X}$ is the set
$\operatorname{UE}(\mathcal{X})=\{(x,y^{\ast}(x)):x\in[0,1]\}.$
For any vectors $q$ and $q^{\prime}$ in the same Euclidean space, let $[q,q^{\prime}]$ denote the line segment connecting $q$ and $q^{\prime}$.
Lemma 3 (Upper envelope)
The upper envelope of $\mathcal{X}$ equals $\left\{(x(q),y(q)):q\in\cup_{k=0}^{N}[q^{k},q^{k+1}]\right\}$.
For $k=0,\ldots,N+1$ let $(x_{k},y_{k})=(x(q^{k}),y(q^{k}))$. Since the map $q\mapsto(x(q),y(q))$ is linear, the upper envelope of $\mathcal{X}$ can alternatively be expressed as $\cup_{k=0}^{N}[(x_{k},y_{k}),(x_{k+1},y_{k+1})]$. Figure 2 illustrates these $N+1$ line segments in an example with $N=4$.
A.5 Proof of Theorem 2
As argued in the main text, it is clear that the sender cannot benefit from communication if $p\leq\underaccent{\bar}{\nu}$, and the sender can benefit from communication if $\underaccent{\bar}{\nu}<p\leq\bar{p}_{N-1}$. Therefore, it suffices to show that the sender cannot benefit from communication if $p>\bar{p}_{N-1}$. For each buying probability vector $q\in Q$ and each price $p$, the receiver’s utility is given by
$u(q,p)=\sum_{j=0}^{N}\pi_{j}q_{j}(v_{j}-p).$
Recall from Section A.4 the definitions of the vectors $\Delta^{0},\ldots,\Delta^{N}$ and the functions $u_{0},\ldots,u_{N}$. We have
$q$ $=q_{0}\Delta^{0}+\sum_{j=1}^{N}q_{j}\left(\Delta^{j}-\frac{j-1}{j}\Delta^{j-1}\right)$
$=\sum_{k=0}^{N-1}\left(q_{k}-\frac{k}{k+1}q_{k+1}\right)\Delta^{k}+q_{N}\Delta^{N}.$
Therefore,
$u(q,p)=\sum_{k=0}^{N-1}\left(q_{k}-\frac{k}{k+1}q_{k+1}\right)u_{k}(p)+q_{N}u_{N}(p).$
For each $q\in Q$, each coefficient on $u_{k}(p)$ in the summation is nonnnegative. Fix a price $p$ satisfying $p>\bar{p}_{N-1}$. By Lemma 2.iiiii, we have $u_{k}(p)<0$ for $k=1,\ldots,N-1$. By direct computation, we have $u_{0}(p)<0$. We consider three cases.
  1. 1.
    $p>\bar{p}_{N}=\bar{\nu}$. The no-communication outcome is $q=0$. By Lemma 2.iii, $u_{N}(p)<0$. Thus, the unique vector $q\in Q$ satisfying $u(q,p)\geq 0$ is $q=0$. Therefore, every equilibrium induces $q=0$.
  2. 2.
    $p=\bar{p}_{N}=\bar{\nu}$. The sender-preferred no-communication outcome is $q=q^{N}$. By Lemma 2.iii, $u_{N}(p)=0$. For each $q\in Q$, we have: $u(q,p)\geq 0$ if and only if $q_{k}=\tfrac{k}{k+1}q_{k+1}$ for all $k=0,\ldots,N-1$, or equivalently, $q=\lambda q^{N}$ for some $\lambda\in[0,1]$. Thus, in every equilibrium, the induced buying probability is weakly smaller than in the sender-preferred no-communication outcome.
  3. 3.
    $p<\bar{p}_{N}=\bar{\nu}$. The no-communication outcome is $q=q^{N}$. By Lemma 2.iii, $u_{N}(p)>0$. For each $q\in Q$, we have: $u(q,p)\geq u_{N}(p)$ if and only if $q_{N}=1$ and $q_{k}=\tfrac{k}{k+1}q_{k+1}$ for all $k=0,\ldots,N-1$. Thus, every equilibrium induces $q=q^{N}$.
A.6 Proof of Theorem 3
First, we introduce notation. For each $k=0,\ldots,N+1$, let $(x_{k},y_{k})=(x(q^{k}),y(q^{k}))$. Observe that
$y_{k}-px_{k}=\sum_{j=0}^{N}\pi_{j}q^{k}_{j}(v_{j}-p).$
In particular, for $k=0,\ldots,N-1$, applying (6) gives
$y_{k}-px_{k}-(y_{k+1}-px_{k+1})=\sum_{j=0}^{N}\pi_{j}(q^{k}_{j}-q^{k+1}_{j})(v_{j}-p)=\frac{u_{k}(p)}{k+1}.$ (8)
Now we turn to the main proof. Fix $k\in\{1,\ldots,N-1\}$. Fix a price $p\in[\bar{p}_{k-1},\bar{p}_{k}]$. We claim that for every point $(x,y)$ on the upper envelope of $\mathcal{X}$, if $x>x_{k}$, then $y-px<y_{k}-px_{k}$. We complete the proof, assuming this claim. Then we verify the claim.
First, we check Pareto efficiency. Suppose for a contradiction that there exists an equilibrium that Pareto dominates the top-$k$ equilibrium. This equilibrium induces a point $(x,y)$ in $\mathcal{X}$ such that $x\geq x_{k}$ and $y-px\geq y_{k}-px_{k}$, where at least one of the inequalities holds strictly. After replacing $y$ with $y^{\ast}(x)$, these inequalities still hold, so we may assume that $(x,y)$ is on the upper envelope of $\mathcal{X}$. Since $y_{k}=y^{\ast}(x_{k})$, we must have $x>x_{k}$; otherwise both inequalities would hold with equality. Thus, the inequality $y-px\geq y_{k}-px_{k}$ contradicts the claim.
Next, we check sender-optimality. Let $p=\bar{p}_{k}$. By Lemma 2.i, we have $y_{k}-px_{k}=y_{N}-px_{N}$. Suppose for a contradiction that there exists an equilibrium that the sender strictly prefers to the top-$k$ equilibrium. This equilibrium induces a point $(x,y)$ in $\mathcal{X}$ such that $x>x_{k}$ and $y-px\geq y_{N}-px_{N}=y_{k}-px_{k}$. After replacing $y$ with $y^{\ast}(x)$, these inequalities still hold, contrary to the claim.
Finally, we prove the claim. If $(x,y)$ is on the upper envelope of $\mathcal{X}$ and $x>x_{k}$, then by Lemma 3, there exists $k^{\ast}$ in $\{1,\ldots,k\}$ and $\lambda$ in $(0,1]$ such that
$(x,y)=(x_{k^{\ast}},y_{k^{\ast}})+\lambda(x_{k^{\ast}-1}-x_{k^{\ast}},y_{k^{\ast}-1}-y_{k^{\ast}}).$
Therefore,
$y-px-(y_{k}-px_{k})$
$=\left[y-px-(y_{k^{\ast}}-px_{k^{\ast}})\right]+\left[(y_{k^{\ast}}-px_{k^{\ast}})-(y_{k}-px_{k})\right]$
$=\lambda\left[y_{k^{\ast}-1}-px_{k^{\ast}-1}-(y_{k^{\ast}}-px_{k^{\ast}})\right]+\sum_{\ell=k^{\ast}}^{k-1}\left[y_{\ell}-px_{\ell}-(y_{\ell+1}-px_{\ell+1})\right]$
$=\frac{\lambda u_{k^{\ast}-1}(p)}{k^{\ast}}+\sum_{\ell=k^{\ast}}^{k-1}\frac{u_{\ell}(p)}{\ell+1}$
$<0,$
where the second equality follows from (8) and the last inequality follows from Lemma 2.ii.
A.7 Proof of Theorem 4
First, we compute an upper bound on the sender’s equilibrium utility, as a function of the price $p$. For each price $p$, let
$V(p)=\max\left\{p\sum_{j=0}^{N}\pi_{j}q_{j}:q\in Q~{}\text{and}~{}\sum_{j=0}^{N}\pi_{j}q_{j}(v_{j}-p)\geq 0\vee\sum_{j=0}^{N}\pi_{j}q_{j}^{N}(v_{j}-p)\right\}.$
Equivalently,
$V(p)=\max\left\{px:(x,y)\in\mathcal{X}~{}\text{and}~{}y-px\geq 0\vee(y_{N}-px_{N})\right\}.$ (9)
Moreover, we can further restrict the maximization in (9) to the upper envelope of $\mathcal{X}$, which is characterized in Lemma 3.
For $k=0,\ldots,N$, let
$\beta_{k}=\frac{y_{k}-y_{k+1}}{x_{k}-x_{k+1}}.$
Since $q^{k}-q^{k+1}$ is proportional to $\Delta^{k}$, we have
$\beta_{k}=\frac{\sum_{j=0}^{N}\pi_{j}\Delta_{j}^{k}v_{j}}{\sum_{j=0}^{N}\pi_{j}\Delta_{j}^{k}}.$
The sequence $\Delta^{0},\ldots,\Delta^{N}$ is strictly increasing in the likelihood ratio order. [^15] Formally, each vector in the sequence can be normalized and then viewed as a probability distribution. The resulting sequence of distributions is strictly increasing in the likelihood ratio order. Since $0\leq v_{0}<\cdots<v_{N}$, we conclude that $0\leq\beta_{0}<\cdots<\beta_{N}$. Therefore, the maximizer in (9) can be computed directly, as follows.
  • If $p\leq\bar{p}_{0}$, then the maximizer is $(x_{0},y_{0})$.
  • If $\bar{p}_{0}<p<\bar{p}_{N-1}$, then the maximizer is the unique point $(x,y)$ on the upper envelope of $\mathcal{X}$ that satisfies $y-px=y_{N}-px_{N}$.
  • If $p=\bar{p}_{N-1}$, then the maximizer is $(x_{N-1},y_{N-1})$.
  • If $\bar{p}_{N-1}<p\leq\bar{p}_{N}$, then the maximizer is $(x_{N},x_{N})$.
  • If $p>\bar{p}_{N}$, then the maximizer is $(x_{N+1},y_{N+1})=(0,0)$.
At $p=\bar{p}_{N-1}$, the function $V$ is upper semicontinuous but not lower semicontinuous. At all other prices $p$, the function $V$ is continuous. Therefore, $V$ has a global maximizer, which must be in $[\bar{p}_{0},\bar{p}_{N}]$. We show that $V$ has a global maximizer in $\{\bar{p}_{0},\ldots,\bar{p}_{N}\}$. It suffices to show that for each $k=0,\ldots,N-1$, the function $V$ is convex over the interval $(\bar{p}_{k},\bar{p}_{k+1})$. Since $V$ is linear over $(\bar{p}_{N-1},\bar{p}_{N})$, it suffices to consider $k=0,\ldots,N-2$.
Fix $k\in\{0,\ldots,N-2\}$ and $p\in(\bar{p}_{k},\bar{p}_{k+1})$. By Lemma 2.i,
$y_{k+1}-px_{k+1}>y_{N}-px_{N}>y_{k}-px_{k},$
so the maximizer must lie on the slope-$\beta_{k}$ line segment between $(x_{k+1},y_{k+1})$ and $(x_{k},y_{k})$. Thus, the maximizer is $(x,y)=(x,y_{k+1}+\beta_{k}(x-x_{k+1}))$ for some $x$ given by
$y_{k+1}+\beta_{k}(x-x_{k+1})-px=y_{N}-px_{N}.$
Solving for $x$ gives
$x=\frac{px_{N}+y_{k+1}-y_{N}-\beta_{k}x_{k+1}}{p-\beta_{k}}.$
Let $\gamma_{k}=y_{k+1}-y_{N}-\beta_{k}x_{k+1}$. We have
$V(p)=\frac{p^{2}x_{N}+p\gamma_{k}}{p-\beta_{k}}.$
Differentiating twice, we get
$V^{\prime}(p)$ $=\frac{p^{2}x_{N}-2px_{N}\beta_{k}-\gamma_{k}\beta_{k}}{(p-\beta_{k})^{2}},$
$V^{\prime\prime}(p)$ $=\frac{2\beta_{k}(x_{N}\beta_{k}+\gamma_{k})}{(p-\beta_{k})^{3}}.$
By Lemma 2.i, we have
$\bar{p}_{k}=\frac{y_{k}-y_{N}}{x_{k}-x_{N}},$
so $\bar{p}_{k}$ is a weighted average of $\beta_{k},\dots,\beta_{N-1}$ with strictly positive weights. Thus, $p>\bar{p}_{k}>\beta_{k}\geq 0$. Therefore, we have $V^{\prime\prime}(p)\geq 0$ if
$0\leq x_{N}\beta_{k}+\gamma_{k}=y_{k+1}-y_{N}-\beta_{k}(x_{k+1}-x_{N}),$
or equivalently,
$\beta_{k}\leq\frac{y_{k+1}-y_{N}}{x_{k+1}-x_{N}}.$
The right side is a weighted average of $\beta_{k+1},\dots,\beta_{N-1}$, so the inequality holds.
A.8 Example for Theorem 4
Consider the simple example with $N=2$ and $v(\theta)=\lvert\theta\rvert$. The price thresholds are given by
$\displaystyle\bar{p}_{0}$ $\displaystyle=\operatorname{\mathbf{E}}\left[\lvert\theta\rvert\mid\theta_{1}=0\right]=\frac{\pi_{1}/2}{\pi_{0}+\pi_{1}/2},$
$\displaystyle\bar{p}_{1}$ $\displaystyle=\operatorname{\mathbf{E}}\left[\lvert\theta\rvert\mid(T^{1},B^{1})=(1,0)\right]=1,$
$\displaystyle\bar{p}_{2}$ $\displaystyle=\operatorname{\mathbf{E}}\left[\lvert\theta\rvert\mid\theta_{1}=1\right]=\frac{\pi_{1}/2+2\pi_{2}}{\pi_{1}/2+\pi_{2}}.$
Therefore,
$\displaystyle V(\bar{p}_{0})$ $\displaystyle=\bar{p}_{0}=\frac{\pi_{1}/2}{\pi_{0}+\pi_{1}/2},$
$\displaystyle V(\bar{p}_{1})$ $\displaystyle=\bar{p}_{1}(\pi_{1}+\pi_{2})=\pi_{1}+\pi_{2},$
$\displaystyle V(\bar{p}_{2})$ $\displaystyle=\bar{p}_{2}(\pi_{1}/2+\pi_{2})=\pi_{1}/2+2\pi_{2}.$
Since $\pi_{0}=1-\pi_{1}-\pi_{2}$, we have $V(\bar{p}_{1})\geq V(\bar{p}_{0})$ if and only if
$\pi_{1}+\pi_{2}\geq\pi_{1}/2+(\pi_{1}/2+\pi_{2})(\pi_{1}+\pi_{2}),$
or equivalently,
$\pi_{1}/2+\pi_{2}\geq(\pi_{1}/2+\pi_{2})(\pi_{1}+\pi_{2}),$
which holds because $\pi_{1}+\pi_{2}\leq 1$. Therefore, either the top-11 11 equilibrium or the top-22 22 equilibrium is optimal. The top-11 11 equilibrium is optimal if $\pi_{1}\geq 2\pi_{2}$ and the top-22 22 equilibrium is optimal if $\pi_{1}\leq 2\pi_{2}$. Note that
$\frac{V(\bar{p}_{1})}{V(\bar{p}_{2})}=\frac{\bar{p}_{1}}{\bar{p}_{2}}\cdot\frac{\pi_{1}+\pi_{2}}{\pi_{1}/2+\pi_{2}}=\frac{\pi_{1}/2+\pi_{2}}{\pi_{1}/2+2\pi_{2}}\cdot\frac{\pi_{1}+\pi_{2}}{\pi_{1}/2+\pi_{2}}.$
So the price ratio and the buying probability ratio are both increasing in $\pi_{1}/\pi_{2}$.
A.9 Proof of Theorem 5
Recall that the favorability order is defined in Section A.1. We begin with two statistical lemmas about this order.
Lemma 4 (Favorability)
Given $n,k\in[N-1]$, the favorability of the event $(T_{n}^{k},B_{n}^{N-k})=(t,b)$ is strictly increasing in $(t,b)$ over the range $S(n,k)$, with respect to the product order.
Lemma 5 (Favorability across equilibria)
For each $n\in[N-1]$, the following hold:
  1. 1.
    The favorability of the event $T_{n}^{k}=0$ is strictly increasing in $k$ over the range $[N]$.
  2. 2.
    The favorability of the event $(T_{n}^{k},B_{n}^{N-k})=(n\wedge k,0)$ is strictly increasing in $k$ over the range $[N]$.
For completeness, we prove both lemmas in Appendix B.
For each $k$, the inequality $\underaccent{\bar}{p}_{k,n}<\bar{p}_{k,n}$ follows from Lemma 4 since $T_{n}^{k}=0$ if and only if $(T_{n}^{k},B_{n}^{N-k})=(0,0)$. The ordering of the price thresholds follows from Lemma 5. It remains to prove the equilibrium condition.
Fix $n,k\in[N-1]$. Under the top-$k$ strategy profile, it is clear that the sender is playing a best response to the receiver’s strategy. We characterize whether the receiver is playing a best response to the sender’s strategy. Suppose that the sender uses the top-$k$ messaging strategy. Further, suppose counterfactually that the receiver could check $n\wedge k$ of the recommended attributes and also $n\wedge(N-k)$ of the unrecommended attributes. By exchangeability, a sufficient statistic for the receiver’s observation is $(T_{n}^{k},B_{n}^{N-k})$. It follows from Lemma 4 that the conditional expectation
$\operatorname{\mathbf{E}}[v(\theta)\mid(T_{n}^{k},B_{n}^{N-k})=(t,b)]$
is strictly increasing in $(t,b)$ over $S(n,k)$, with respect to the product order. Thus, the receiver has a best response that takes one of the following forms: (i) the receiver never buys, or (ii) the receiver buys if and only if $(T_{n}^{k},B_{n}^{N-k})\geq(t^{\ast},b^{\ast})$ for some $(t^{\ast},b^{\ast})\in S(n,k)$. We now separate into cases.
If $\underaccent{\bar}{p}_{k,n}\leq p\leq\bar{p}_{k,n}$, then in the counterfactual game, the receiver has a best response of the form (ii) for some threshold $(t^{\ast},b^{\ast})$ with $t^{\ast}>0$ and $b^{\ast}=0$. Equivalently, the receiver buys if and only if $T_{n}^{k}\geq t^{\ast}$. This buying rule can be implemented in the original game by the receiver’s top-$k$ strategy. Thus, the receiver’s top-$k$ strategy is a best response in the original game (since the set of deviations is smaller).
If $p<\underaccent{\bar}{p}_{k,n}$, then the receiver’s unique best response in the counterfactual game is to always buy. In the original game, this strategy is still feasible, and hence constitutes a strictly profitable deviation from the top-$k$ strategy.
If $p>\bar{p}_{k,n}$, then the receiver has a best response in the counterfactual game either of the form (i), or of the form (ii) for some threshold $(n\wedge k,b^{\ast})$ with $b^{\ast}>0$. Each of these buying rules can be implemented in the original game. On the other hand, the receiver’s top-$k$ strategy is not a best response in the counterfactual game (because the receiver buys with positive probability when $B_{n}^{N-k}=0$). We conclude that the receiver’s top-$k$ strategy is not a best response in the original game.
A.10 Proof of Lemma 2
  1. (i)
    We first consider the main case with $k>0$. Fix $k\in\{1,\ldots,N-1\}$. For $j=0,\ldots,N$, observe that
    $q_{j}^{k}-q_{j}^{N}$ $=\frac{j(N-k)}{kN}[j<k]+\frac{N-j}{N}[j\geq k]$
    $=\frac{N-k}{N}\cdot\mathbf{P}\left((T^{k},B^{N-k})=(1,0)\mid\lvert\theta\rvert=j\right).$
    Therefore,
    $\sum_{j=0}^{N}\pi_{j}(q_{j}^{k}-q_{j}^{N})(v_{j}-p)$
    $=\frac{N-k}{N}\sum_{j=0}^{N}\pi_{j}\mathbf{P}\left((T^{k},B^{N-k})=(1,0)\mid\lvert\theta\rvert=j\right)(v_{j}-p)$
    $=\frac{N-k}{N}\operatorname{\mathbf{E}}\left\{(v(\theta)-p)\left[(T^{k},B^{N-k})=(1,0)\right]\right\}$
    $=\frac{N-k}{N}\mathbf{P}\left((T^{k},B^{N-k})=(1,0)\right)\left(\bar{p}_{k}-p\right).$
    The last line is nonnegative if and only if $p\leq\bar{p}_{k}$.
    Now we consider the edge case $k=0$. Since $q^{0}=\bm{1}$, we have
    $\sum_{j=0}^{N}\pi_{j}(q_{j}^{0}-q_{j}^{N})(v_{j}-p)$ $=\sum_{j=0}^{N}\pi_{j}\frac{N-j}{N}(v_{j}-p)$
    $=\sum_{j=0}^{N}\pi_{j}\mathbf{P}(\theta_{1}=0\mid\lvert\theta\rvert=j)(v_{j}-p)$
    $=\operatorname{\mathbf{E}}\left\{(v(\theta)-p)[\theta_{1}=0]\right\}$
    $=\mathbf{P}(\theta_{1}=0)(\bar{p}_{0}-p).$
    The last line is nonnegative if and only if $p\leq\bar{p}_{k}$.
  2. (ii)
    Fix $k\in\{1,\ldots,N-2\}$. Since the function $u_{k}$ is strictly decreasing, it suffices to prove that $u_{k}(\bar{p}_{k})<0$. Applying (i), for $k$ and $k+1$, with $p=\bar{p}_{k}$, we conclude that
    00$\displaystyle$ 0 $\displaystyle=\sum_{j=0}^{N}\pi_{j}(q_{j}^{k}-q_{j}^{N})(v_{j}-\bar{p}_{k}),$
    00$\displaystyle$ 0 $\displaystyle<\sum_{j=0}^{N}\pi_{j}(q_{j}^{k+1}-q_{j}^{N})(v_{j}-\bar{p}_{k}).$
    Subtracting the second summation from the first, we get
    $0>\sum_{j=0}^{N}\pi_{j}(q_{j}^{k}-q_{j}^{k+1})(v_{j}-\bar{p}_{k})=\frac{u_{k}(\bar{p}_{k})}{k+1}.$
  3. (iii)
    For $k=N-1$, the result is immediate from (i) since
    $\sum_{j=0}^{N}\pi_{j}(q_{j}^{N-1}-q_{j}^{N})(v_{j}-p)=\frac{u_{N-1}(p)}{N}.$
    For $k=N$, we have
    $u_{N}(p)$ $=\sum_{j=0}^{N}\pi_{j}(j/N)(v_{j}-p)$
    $=\sum_{j=0}^{N}\pi_{j}\mathbf{P}(\theta_{1}=1\mid\lvert\theta\rvert=j)(v_{j}-p)$
    $=\operatorname{\mathbf{E}}\left\{(v(\theta)-p)[\theta_{1}=1]\right\}$
    $=\mathbf{P}(\theta_{1}=1)(\bar{p}_{N}-p).$
    That last line is nonnegative if and only if $p\leq\bar{p}_{N}$.
A.11 Proof of Lemma 3
Let $Q^{\prime}$ be the set of all vectors $q\in Q$ for which there exists $k\in\{0,\ldots,N-1\}$ such that $q_{k}>\tfrac{k}{k+1}q_{k+1}$ and $q_{k+1}<1$. We claim that $\cup_{k=0}^{N}[q^{k},q^{k+1}]=Q\setminus Q^{\prime}$. First, we check that $\cup_{k=0}^{N}[q^{k},q^{k+1}]\subset Q\setminus Q^{\prime}$. For each $k=0,\ldots,N$, if $q$ is in $[q^{k},q^{k+1}]$, then $q_{j}=1$ for all $j\geq k+1$, and $q_{j}=\tfrac{j}{j+1}q_{j+1}$ for all $j\leq k-1$. Thus, $q$ is not in $Q^{\prime}$. Next, we check that $Q\setminus Q^{\prime}\subset\cup_{k=0}^{N}[q^{k},q^{k+1}]$. Fix $q\in Q\setminus Q^{\prime}$. If $q=\bm{1}$, then $q$ is in $[q^{0},q^{1}]$, so we may assume that $q\neq\bm{1}$. Let $k$ be the largest index $j$ such that $q_{j}<1$. If $k<N$, then $q_{j}=1$ for all $j>k$, and $q_{k}\geq\tfrac{k}{k+1}q_{k+1}$. Applying the definition of $Q^{\prime}$ repeatedly, we conclude that $q_{j}=(j/k)q_{k}$ for all $j=k-1,k-2,\ldots,0$. Therefore, $q$ is in $[q^{k},q^{k+1}]$.
Let $\mathcal{X}_{0}=\{(x(q),y(q)):q\in Q\setminus Q^{\prime}\}$. We prove that $\operatorname{UE}(\mathcal{X})=\mathcal{X}_{0}$. Since the sets $\operatorname{UE}(\mathcal{X})$ and $\mathcal{X}_{0}$ can each expressed as the graph of a function on $[0,1]$, it suffices to show that $\operatorname{UE}(\mathcal{X})\subseteq\mathcal{X}_{0}$. To prove this inclusion, we show that $(x(q),y(q))\notin\operatorname{UE}(\mathcal{X})$ for each $q\in Q^{\prime}$. Fix $q\in Q^{\prime}$. By the definition of $Q^{\prime}$, there exists $k\in\{0,\ldots,N-1\}$ such that $q_{k}>\tfrac{k}{k+1}q_{k+1}$ and $q_{k+1}<1$. We perturb the vector $q$ by scaling down $q_{0},\ldots,q_{k}$ and increasing $q_{k+1}$ in such a way that the expected buying probability is preserved. Given $\varepsilon>0$, define $\tilde{q}=(\tilde{q}_{0},\ldots,\tilde{q}_{N})$ by
$\tilde{q}_{j}=\begin{cases}(1-\varepsilon)q_{j}&\text{if}~{}j\leq k,\\ q_{j}+\varepsilon\sum_{\ell=0}^{k}(\pi_{\ell}/\pi_{j})q_{\ell}&\text{if}~{}j=k+1,\\ q_{j}&\text{if}~{}j>k+1.\end{cases}$
If $\varepsilon$ is chosen sufficiently small, then $\tilde{q}_{k}>\tfrac{k}{k+1}\tilde{q}_{k+1}$ and $\tilde{q}_{k+1}<1$. In this case, the vector $\tilde{q}$ is in $Q$. By construction, $x(\tilde{q})=x(q)$ and $y(\tilde{q})>y(q)$. Therefore, $(x(q),y(q))$ is not in $\operatorname{UE}(\mathcal{X})$.
Appendix B Supplemental appendix
We assume that $\wedge$ is performed before addition and subtraction. For example, $a-b\wedge c$ means $a-(b\wedge c)$.
B.1 Proof of Lemma 4
Fix $n,k\in[N-1]$. For each $(t,b)\in S(n,k)$ and $0\leq j\leq N$, let
$f(t,b|j)=\mathbf{P}\left((T_{n}^{k},B_{n}^{N-k})=(t,b)\mid\lvert\theta\rvert=j\right).$
We separate into cases. For $0\leq t<n\wedge k$, we have
$f(t,0|j)=\begin{cases}\frac{\binom{j\wedge k}{t}\binom{k-j\wedge k}{n\wedge k-t}\binom{(N-j)\wedge(N-k)}{n\wedge(N-k)}}{\binom{k}{n\wedge k}\binom{N-k}{n\wedge(N-k)}}&\text{if}~{}t\leq j\leq k-n\wedge k+t,\\ 0&\text{otherwise}.\end{cases}$
And
$f(n\wedge k,0|j)=\begin{cases}\frac{\binom{j\wedge k}{n\wedge k}\binom{(N-j)\wedge(N-k)}{n\wedge(N-k)}}{\binom{k}{n\wedge k}\binom{N-k}{n\wedge(N-k)}}&\text{if}~{}n\wedge k\leq j\leq N-n\wedge(N-k),\\ 0&\text{otherwise}.\end{cases}$
For $1\leq b\leq n\wedge(N-k)$, we have
$f(n\wedge k,b|j)=\begin{cases}\frac{\binom{j-k}{b}\binom{(N-j)\wedge(N-k)}{n\wedge(N-k)-b}}{\binom{N-k}{n\wedge(N-k)}}&\text{if}~{}k+b\leq j\leq N-n\wedge(N-k)+b,\\ 0&\text{otherwise}.\end{cases}$
Note that the supports are strictly increasing (with respect to the strong set order) in $(t,b)$.
Thus, for $1\displaystyle 0\leq t\leq n\wedge k-$ 1 and $t+1\leq j\leq k-n\wedge k+t$, we have
$\frac{f(t+1,0|j)}{f(t,0|j)}$ $=\frac{\binom{j\wedge k}{t+1}\binom{k-j\wedge k}{n\wedge k-t-1}}{\binom{j\wedge k}{t}\binom{k-j\wedge k}{n\wedge k-t}}$
$=\frac{j\wedge k-t}{t+1}\cdot\frac{n\wedge k-t}{k-j\wedge k-n\wedge k+t+1}.$
This ratio is weakly increasing in $j$. For $1\displaystyle 0\leq b\leq n\wedge(N-k)-$ 1 and $k+b+1\leq j\leq N-n\wedge(N-k)+b$, we have
$\frac{f(n\wedge k,b+1|j)}{f(n\wedge k,b|j)}$ $=\frac{\binom{j-k}{b+1}\binom{(N-j)\wedge(N-k)}{n\wedge(N-k)-b-1}}{\binom{j-k}{b}\binom{(N-j)\wedge(N-k)}{n\wedge(N-k)-b}}$
$=\frac{j-k-b}{b+1}\cdot\frac{n\wedge(N-k)-b}{(N-j)\wedge(N-k)-n\wedge(N-k)+b+1}.$
This ratio is weakly increasing in $j$.
B.2 Proof of Lemma 5
Fix $n\in[N-1]$.
  1. 1.
    For $1\leq k\leq N$ and $0\leq j\leq N$, let
    $f(k|j)=\mathbf{P}(T_{n}^{k}=0\mid\lvert\theta\rvert=j)=\begin{cases}\frac{\binom{k-j}{n\wedge k}}{\binom{k}{n\wedge k}}&\text{if}~{}j\leq k-n\wedge k,\\ 0&\text{otherwise}.\end{cases}$
    Note that the supports are weakly increasing (with respect to the strong set order) in $k$, strictly so if $k\geq n$.
    Thus, for $1\leq k\leq N$ and $1\displaystyle 0\leq j\leq k-n\wedge k-$ 1, we have
    $\frac{f(k|j+1)}{f(k|j)}=\frac{\binom{k-j-1}{n\wedge k}}{\binom{k-j}{n\wedge k}}=\frac{k-j-n\wedge k}{k-j}.$
    This ratio is weakly increasing in $k$, strictly so if $k\leq n$.
  2. 2.
    For this part, we use the convention that $B_{n}^{0}=0$ and $\tbinom{0}{0}=1$. For $1\leq k\leq N$ and $0\leq j\leq N$, let
    $f(k|j)$ $=\mathbf{P}\left((T_{n}^{k},B_{n}^{N-k})=(n\wedge k,0)\mid\lvert\theta\rvert=j\right)$
    =$=\begin{cases}\frac{\binom{j\wedge k}{n\wedge k}\binom{(N-j)\wedge(N-k)}{n\wedge(N-k)}}{\binom{k}{n\wedge k}\binom{N-k}{n\wedge(N-k)}}&\text{if}~{}n\wedge k\leq j\leq N-n\wedge(N-k),\\ 0&\text{otherwise}.\end{cases}$
    Note that the supports are weakly increasing (with respect to the strong set order) in $k$, strictly so if $k\leq n$ or $k\geq N-n$.
    Thus, for $1\leq k\leq N$ and $n\wedge k\leq j\leq N-n\wedge(N-k)-1$, we have
    $\frac{f(k|j+1)}{f(k|j)}$ $=\frac{\binom{(j+1)\wedge k}{n\wedge k}\binom{(N-j-1)\wedge(N-k)}{n\wedge(N-k)}}{\binom{j\wedge k}{n\wedge k}\binom{(N-j)\wedge(N-k)}{n\wedge(N-k)}}$
[…]
    This expression is weakly increasing in $k$. Moreover, if $n<N-n$, then for $n\leq k\leq N-n-1$, we have
    $\frac{f(k+1|k+1)f(k|k)}{f(k+1|k)f(k|k+1)}>1.$